16-4(y+1)=y+6=5y
2/3(a-6)=2(a+4)
-35-5(x-3)=x+2(x-2)

Answer by checkley71(8403)

16-4(Y-1)=Y+6=5Y HERE YOU HAVE 1 TOO MANY = SIGNS.
--------------------------------------------
2/3(A-6=2(A+4)
(2A-12/3=2A+8 CROSS MULTIPLY
3(2A+8)=2A-12
6A+24=2A-12
6A-2A=-12-24
4A=-36
A=-36/4
A=-9 ANSWER.
-----------------------------------------------
-35-5(X-3)=X+2(X-2)
-35-5X+15=X+2X-4
-5X-X-2X=35-15-4
-8X=16
X=16/-8
X=-2 ANSWER.